2x^2-3x=-5
Use quadratic formula to solve equation

Answer by mananth(16949)

2x^2-3x=-5
+5 +5
2x^2-3x+5=0
...
discriminant = -31
...
x1= ((-(-3)+sqrt(-31))/4
x1= 3/4 +sqrt(-31)/4
x1= 0.75 + 1.3919410907075053i
...
x2= ((-(-3)-sqrt(-31))/4
x2= 3/4 -sqrt(-31)/4
x2= 0.75 - 1.3919410907075053i
...
m.ananth@hotmail.ca